OpenStudy (anonymous):

ALGEBRA: The perimeter of a rectangle is 96cm. The longest side is 3 times as long as the shorter side. Form an equation to find the length of the longest side of the rectangle. - How do I figure out this and similar question.. steps please!

OpenStudy (id21):

x- width 3x- length \[2\times(x+3x)=96 cm\] - perimeter

OpenStudy (id21):

\[2x+6x=96; 8x=96; x=12cm\]-tkis is width the length is \[3x=3 \times 12 = 36cm\]